11. Let x = ln 3 and y = ln 5. Write the following expressions in terms of x and y.

a. ln(5/3)
b. ln45

Answer by Alan3354(69443)

Let x = ln 3 and y = ln 5. Write the following expressions in terms of x and y.
------------
Log(a/b) = log(a) - log(b)
log(a*b) = log(a) + log(b)
====================
a. ln(5/3)
5/3 = 5/3
=======
b. ln45
45 = 3*3*5